# Effective action for $\phi^{4}$-Yukawa theory via 2PI formalism in the inflationary de Sitter spacetime

### v2

@@ 4 Conclusions
A possible caveat to our analysis, as we mentioned earlier as well, is the estimation of the value of the secular logarithm (cf., discussion below 54). As we have discussed, we could not find any formal or systematic way to make resummation, if any, for such terms.
These results can be applied or extended in different directions. For example, what should be the scalar correlation functions with the effective potential of 71? How do they differ quantitatively and qualitatively from that of the simple $\lambda\phi^{4}$ potential? Inclusion of gauge fields would also be interesting. Finally and most importantly, we have focussed only on the local part of the effective potential or part of the propagator relevant for that. How do we compute the non-local parts? This involves solving the Schwinger-Dyson equations with non-local self energy contributions, and we need to use the Schwinger-Keldysh formalism. For scalar field theory some asymptotic analysis can be seen in [43, 51] with non-local self energy. All these seem to be challenging as well as important tasks and we hope to come back to at least some of them in the near future.
@@ Acknowledgements
KR’s research is partly supported by the fellowship from Council of Scientific and Industrial Research, Government of India (File No. 09/096(0987)/2019-EMR-I).
@@ Appendix A Three loop 2PI vacuum diagram at ${\cal O}(\lambda^{2})$
In order to introduce the in-in formalism, we write from 74 for the anti-time ordering
@@
From 74, 75 and using 76, we have the in-in matrix elements
@@
$i\Delta^{4}(x,x^{\prime})=A^{4}+4A^{3}B+6A^{2}B^{2}+4AB^{3}+B^{4}+4A^{3}C$ (80)
The rest of the terms containing $C(x,x^{\prime})$ are finite and drop out as $\epsilon\to 0$. We have
@@
Thus from 82, we have
@@
We wish to first compute below the local contribution in 79, arising from the Feynman propagator (+++++ +). This will require renormalisation, following which we will obtain finite local contribution. After this, we shall compute the aforementioned deep infrared, non-local contribution separately. Thus for $i\Delta^{4}_{++}(x,x^{\prime})$ in 79, the contribution from the $A_{++}^{4}$ term, 81, reads after using 83,
@@
We next consider the $4A_{++}^{3}B_{++}$ term in 79, 81. Using the second and the third of 83, we have for the local contribution
@@
Let us now consider the $6A^{2}B^{2}$ term in 81. Using the last three equations of 83, we have for the local contribution
@@
It can be shown that the $AB^{3}$ and $B^{4}$ terms do not contribute to any local terms in the effective action. The only remaining term relevant for this purpose will be the $4A^{3}C$ term in 81,
@@
where the sgn stands for the sign of $(\eta-\eta^{\prime})$ and it is zero for $\eta=\eta^{\prime}$. Putting things together now, 87 becomes
@@
It is easy to check by expanding the Gamma functions within parenthesis that the above term is not divergent. Thus we may set $d=4$, $a^{2\epsilon}=1$ above and hence the whole term can be absorbed in a finite cosmological constant counterterm, $\delta\Lambda/8\pi G$. Combining now 84, 85, 86 and 89, we have the three loop, local contribution
@@
$-\frac{\delta\lambda}{4!}\int a^{d}d^{d}x\langle\phi^{4}\rangle=-\frac{\delta\lambda}{2^{3}}\int a^{d}d^{d}x\ i\Delta^{2}(x,x)=-\frac{\delta\lambda H^{4-2\epsilon}\Gamma^{2}(2-\epsilon)}{2^{7-2\epsilon}\pi^{4-\epsilon}\Gamma^{2}(1-\epsilon/2)}\int a^{d}d^{d}x\left(\frac{1}{\epsilon^{2}}+\frac{2\ln a}{\epsilon}+\ln^{2}a\right)$
@@
$-\frac{\delta m^{2}}{2}\int a^{d}d^{d}x\langle\phi^{2}\rangle=-\frac{\delta m^{2}H^{2-\epsilon}\Gamma(2-\epsilon)}{2^{3-\epsilon}\pi^{2-\epsilon/2}\Gamma(1-\epsilon/2)}\int a^{d}d^{d}x\left(\ln a+\frac{1}{\epsilon}\right)$ (92)
@@
Let us now come to the non-local part of $\Gamma_{2}^{3-{\rm loop}}$. This will yield the leading deep infrared contribution of ${\cal O}(\ln^{4}a)$. We shall use the in-in formalism to do this. Now, instead of using the exact expressions of 83, we wish to use the IR effective formalism described in e.g. [50]. We have from 79 appropriate for the in-in and non-local contribution
@@
We now rewrite 95 in the spatial momentum space. In the deep IR, super-Hubble limit, we have $H\lesssim k\lesssim Ha$, where $k=|\vec{k}|$.
@@
×$\times\left(i\Delta_{-+}(k_{3},\eta,\eta^{\prime})i\Delta_{-+}(|{\vec{k}_{1}}+{\vec{k}_{2}}+{\vec{k}_{3}}|,\eta,\eta^{\prime})+{\rm c.c.}\right)\Theta(Ha^{\prime}-k_{1})\Theta(Ha^{\prime}-k_{2})\Theta(Ha^{\prime}-k_{3})\Theta(Ha^{\prime}-|{\vec{k}_{1}}+{\vec{k}_{2}}+{\vec{k}_{3}}|)$
@@
96, combined with 94 gives the renormalised, leading late time expression for $i\Gamma_{2}^{3-{\rm loop}}$ at ${\cal O}(\lambda^{2})$, yielding 38.
Appendix B Massive propagators and their coincidence limits
@@
If we multiply 108 by the fermion mass $M$, the corresponding quantity represents the trace of the free fermion’s energy-momentum tensor. Accordingly, it can be renormalised via a cosmological constant counterterm, giving
@@ Appendix C Local self energy integral with fermion propagators for 56
where the fermion propagator is given by 108. The gamma matrices appearing above are flat space, satisfying $[\gamma^{\mu},\gamma^{\nu}]_{+}=-2\eta^{\mu\nu}\ {\bf I}_{d\times d}$, appropriate for our mostly positive metric signature. We have also used 83 in the second equality.
For the two loop Yukawa vacuum graph, we need to put $x=x^{\prime}$ in 111, where the coincidence scalar propagator $iG(x,x)$ is given by 104. Also, while using in the main text, the mass terms appearing above needs to replaced by the respective effective dynamical masses.
@@ Appendix D On the difference of results between 2PI renormalisation and the standard 1PI perturbative method
Using now 8 and 83, one can see that the square of the Feynman propagator generates a secular logarithm, and by choosing then $\delta\lambda_{\rm fin}$ appropriately, we can remove the aforementioned divergence. We refer our reader to [65] and references therein for the standard two loop renormalisation of the $\phi^{4}$-Yukawa theory in de Sitter. Note in particular that the finite secular logarithm term generated by the Yukawa vacuum loop is quadratic, opposed to the non-perturbative result of 69.

### v3

@@ 4 Conclusions
Let us now come to the issue of possible phenomenological implications of our effective potential, 71. Following [66], we may treat the background field $v$ to be stochastic at late times, and compute the inflationary correlation functions. By computing the two or higher point correlators in the spatial momentum space and the associated power spectrum and non-Gaussianity, we may reasonably hope to put some bound on the Yukawa coupling, for example, via the observed constraint on the spectral index, $1-n_{s}\lesssim 0.04$. As a warm up, let us consider here computing the stochastic average $\langle v^{2}(x)\rangle$ [54],
$\langle v^{2}(x)\rangle_{\rm stoch.}=N^{-1}\int_{-\infty}^{\infty}dv\ v^{2}e^{-8\pi V_{\rm eff}(v)/3}$
where $N=\int_{-\infty}^{\infty}dv\ e^{-8\pi V_{\rm eff}(v)/3}$ is the normalisation. The above has been plotted in the second of 10. Since the fermion mass is proportional to the Yukawa coupling, $M_{\rm dyn.eff}=gv$, the plot qualitatively shows the possible effect of the fermion effective mass onto the two point correlation function. Perhaps this might be an interesting result indicating that the Yukawa coupling makes the corresponding power spectrum blue tilted. However, a more realistic such computation should involve a quasi-de Sitter metric and associated slow roll parameters. In this case, we must treat the scalar field as the inflaton.
Another interesting thing to look into would be the post inflationary reheating phenomenon with this effective potential. The dynamical mass residing in the effective potential contains the inflationary Hubble rate. At the end of the inflation, this Hubble rate would decay reducing the value of the dynamical mass and thereby releasing a huge amount of rest mass energy. Could this energy reheat (or at least contribute to some extent in reheating) the universe? Certainly, this would require a coupling between the scalar and the photons. One can also look into the $O(N)$ symmetric scalar field theory without the Yukawa interaction for this purpose. Finally we note that, we have focussed only on the local part of the effective potential. How do we compute the non-local parts? This involves solving the Schwinger-Dyson equations with non-local self energy contributions, and we need to use the Schwinger-Keldysh formalism. For scalar field theory some asymptotic analysis can be seen in [43, 51] with the two loop non-local self energy. Most importantly, how does this non-local part affect the correlation functions and the phenomenology of the spectral index? Can this part affect the small scale scenario like the Jean’s instability during a gravitational collapse? All these seem to be challenging as well as important tasks, and we hope to come back to at least some of them in the near future.
@@ Acknowledgements
KR’s research was partly supported by the fellowship from Council of Scientific and Industrial Research, Government of India (File No. 09/096(0987)/2019-EMR-I). The authors would like to sincerely acknowledge anonymous referee for many useful comments and suggestions.
@@ Appendix A Three loop 2PI vacuum diagram at ${\cal O}(\lambda^{2})$
In order to introduce the in-in formalism, we write from 76 for the anti-time ordering
@@
From 76, 77 and using 78, we have the in-in matrix elements
@@
$i\Delta^{4}(x,x^{\prime})=A^{4}+4A^{3}B+6A^{2}B^{2}+4AB^{3}+B^{4}+4A^{3}C,$ (82)
Note that we have omitted some terms containing $C(x,x^{\prime})$. This is because the same is vanishing for either $\epsilon\to 0$ or $\Delta x^{2}\to 0$, or both, 8. Also, the term $B(x,x^{\prime})$ is finite as $\epsilon\to 0$. Thus first of all, all terms containing only $B$ and $C$ drop out. Let us now consider the term $A^{2}BC$. We can see from 84, 85 below that this term behaves as $\sim\delta^{d}(x-x^{\prime})B(x,x^{\prime})C(x,x^{\prime})/\epsilon$, which vanishes in the coincidence limit due to the $\delta$-function, while integrating in 81. Likewise all the other terms containing A,BA,\ B and $C$ vanish. It was pointed out earlier in [18] that in order to have any non-vanishing contribution associated with $C(x,x^{\prime})$, it must be multiplied with cubic or higher power of $A$. We will show below that the $A^{3}C$ term appearing in 82 contains no UV divergence.
We next compute
@@
Thus from 84, we have
@@
We wish to first compute below the local contribution in 81, arising from the Feynman propagator (++++). This will require renormalisation, following which we will obtain finite local contribution. After this, we shall compute the aforementioned deep infrared, non-local contribution separately. Thus for $i\Delta^{4}_{++}(x,x^{\prime})$ in 81, the contribution from the $A_{++}^{4}$ term, 83, reads after using 85,
@@
We next consider the $4A_{++}^{3}B_{++}$ term in 81, 83. Using the second and the third of 85, we have for the local contribution
@@
Let us now consider the $6A^{2}B^{2}$ term in 83. Using the last three equations of 85, we have for the local contribution
@@
It can be shown that the $AB^{3}$ and $B^{4}$ terms do not contribute to any local terms in the effective action. The only remaining term relevant for this purpose will be the $4A^{3}C$ term in 83,
@@
where the sgn stands for the sign of $(\eta-\eta^{\prime})$ and it is zero for $\eta=\eta^{\prime}$. Putting things together now, 89 becomes
@@
It is easy to check by expanding the Gamma functions within parenthesis that the above term is not divergent. Thus we may set $d=4$, $a^{2\epsilon}=1$ above and hence the whole term can be absorbed in a finite cosmological constant counterterm, $\delta\Lambda/8\pi G$. Combining now 86, 87, 88 and 91, we have the three loop, local contribution
@@
$\frac{\delta\lambda}{4!}\int a^{d}d^{d}x\langle\phi^{4}\rangle=-\frac{\delta\lambda}{2^{3}}\int a^{d}d^{d}x\ i\Delta^{2}(x,x)=-\frac{\delta\lambda H^{4-2\epsilon}\Gamma^{2}(2-\epsilon)}{2^{7-2\epsilon}\pi^{4-\epsilon}\Gamma^{2}(1-\epsilon/2)}\int a^{d}d^{d}x\left(\frac{1}{\epsilon^{2}}+\frac{2\ln a}{\epsilon}+\ln^{2}a\right)$
@@
$\frac{\delta m^{2}}{2}\int a^{d}d^{d}x\langle\phi^{2}\rangle=-\frac{\delta m^{2}H^{2-\epsilon}\Gamma(2-\epsilon)}{2^{3-\epsilon}\pi^{2-\epsilon/2}\Gamma(1-\epsilon/2)}\int a^{d}d^{d}x\left(\ln a+\frac{1}{\epsilon}\right)$ (94)
@@
Let us now come to the non-local part of $\Gamma_{2}^{3-{\rm loop}}$. This will yield the leading deep infrared contribution of ${\cal O}(\ln^{4}a)$. We shall use the in-in formalism to do this. Now, instead of using the exact expressions of 85, we wish to use the IR effective formalism described in e.g. [50]. We have from 81 appropriate for the in-in and non-local contribution
@@
We now rewrite 97 in the spatial momentum space. In the deep IR, super-Hubble limit, we have $H\lesssim k\lesssim Ha$, where $k=|\vec{k}|$.
@@
$\times\left(i\Delta_{-+}(k_{3},\eta,\eta^{\prime})i\Delta_{-+}(|{\vec{k}_{1}}+{\vec{k}_{2}}+{\vec{k}_{3}}|,\eta,\eta^{\prime})+{\rm c.c.}\right)\Theta(Ha^{\prime}-k_{1})\Theta(Ha^{\prime}-k_{2})\Theta(Ha^{\prime}-k_{3})\Theta(Ha^{\prime}-|{\vec{k}_{1}}+{\vec{k}_{2}}+{\vec{k}_{3}}|)$
@@
98, combined with 96 gives the renormalised, leading late time expression for $i\Gamma_{2}^{3-{\rm loop}}$ at ${\cal O}(\lambda^{2})$, yielding 38.
Appendix B Propagators and their coincidence limits
@@
If we multiply 110 by the fermion mass $M$, the corresponding quantity represents the trace of the free fermion’s energy-momentum tensor. Accordingly, it can be renormalised via a cosmological constant counterterm, giving
@@ Appendix C Local self energy integral with fermion propagators for 56
where the fermion propagator is given by 110. The gamma matrices appearing above are flat space, satisfying $\gamma^{\mu},\gamma^{\nu}]_{+}=-2\eta^{\mu\nu}\ {\bf I}_{d\times d}$, appropriate for our mostly positive metric signature. We have also used 85 in the second equality.
For the two loop Yukawa vacuum graph, we need to put $x=x^{\prime}$ in 113, where the coincidence scalar propagator $iG(x,x)$ is given by 106. Also, while using in the main text, the mass terms appearing above needs to replaced by the respective effective dynamical masses.
@@ Appendix D On the difference of results between 2PI renormalisation and the standard 1PI perturbative method
Using now 8 and 85, one can see that the square of the Feynman propagator generates a secular logarithm, and by choosing then $\delta\lambda_{\rm fin}$ appropriately, we can remove the aforementioned divergence. We refer our reader to [65] and references therein for the standard two loop renormalisation of the $\phi^{4}$-Yukawa theory in de Sitter. Note in particular that the finite secular logarithm term generated by the Yukawa vacuum loop is quadratic, opposed to the non-perturbative result of 69.